13th October 2018

Media storage currently under investigation

Earlier this morning a major temperature spike has been detected in the data center. This has caused several problems which we are identifying right now. While our main media storage is not accessible right now, our back-up storage is accessible and was last synced tonight at 01:00.

We are working on a plan to get the media available as soon as possible.

This issue will be updated no later than 12:00 CEST.